# Retinotopic Mapping Enhances the Robustness of Convolutional Neural Networks

### v1

@@ 3.1 Training on Transformed Images
We fine-tuned the pre-trained Vgg16 and ResNet101 networks on the ImageNet dataset with retinotopic transformed images. We compared the performance of the networks trained on the original dataset (with a Cartesian mapping) with the performance of our retrained network on the transformed dataset defined in retinotopic space. We found that the networks retrained on the transformed images had similar categorization accuracy as the network pre-trained on regular images (see Figure 1-A & B). This is surprising given that the networks used in this retraining process were previously trained on regular images and that images with a log-polar transformation show a high degree of distortion, in particular a compression of visual information around the fixation point and a degradation of textures in the periphery, see Figure 2. One hypothesis is that the degradation of texture during the frame of reference change may cause the network to rely on shape rather than texture. However, it remains to be investigated whether there is a qualitative change in the features underlying the categorization.
3.2 Robustness to rotations
We then tested the robustness of the networks to rotation. We applied a rotation to the input image around the fixation point with an angle ranging from $-180^{\circ}$ to $+180^{\circ}$ in steps of $15^{\circ}$. We tested each network (Vgg16 or ResNet101) either with raw images or with retinotopic mapping (i.e., Cartesian or retinotopic). We first observed that Vgg16 has a degraded overall performance compared to ResNet101 and that for both networks, image rotation has a detrimental effect on categorization performance, yet less pronounced for ResNet101. We also observed a symmetry of results corresponding to horizontal flip invariance in images, such that we only show results for half of rotation angles.
@@ 3.2 Robustness to rotations
Noteworthy, these results show that while the Vgg16 network retrained and tested on regular images showed some degradation for different rotations, the categorization results were much more invariant for the network including a retinotopic mapping (see Figure 1-C). This phenomenon is a consequence of the translation invariance imposed by the structure of the CNNs. Applied to retinotopic mapping, this translation invariance in retinotopic space is translated into rotation and zoom invariance in visual space (Araujo and Dias, 1997). The performance of ResNet101 with either Cartesian or retinotopic mapping is similar. Surprisingly, although this network was not designed a priori for retinotopic images, we observe a slight but consistent advantage for retinotopic mapping in this retrained network.
3.3 Visual Object Localization
@@
The CNNs described above are designed to categorize images by providing a likelihood value for each of the $1000$ labels. This likelihood is a probability (i.e., a scalar between 00 and $1$) that predicts the probability that the label is present in the image. This allows us to make a binary decision ("present" or "not present"), e.g. by selecting the label corresponding to the highest likelihood. In our setting, we can also take different views of a large image and compute the likelihood for each of them, allowing us to compare which view provides the best likelihood (“bootstrapping”). Views can consist, for example, of focusing on regions of the image centered on different fixation points, with the fixation points aligned on a regular grid in visual space.
@@
This sample is then transformed or not by the retinotopic mapping before being used as input for the corresponding network, see Appendix Figure 6-B & D. Conveniently, a collection of samples for different fixation points can be processed as a single batch. This protocol defines a likelihood map for any given network as the likelihood of categorizing the presence of a label of interest (here "an animal") inferred at regularly spaced fixation points in the image. A first qualitative exemple was process using $224\times 224$ fixation points, and a 30% ratio, see Figure 3. As can be observed qualitatively in the example image, the Cartesian representation induces uniform activation across the whole image, while the retinotopic mapping induces differential activation inside and outside the object of interest.
Figure 3: Likelihood maps computed on a representative image using $224\times 224$ points of fixation, (top) with a Vgg16, or (bottom) a ResNet101) network, trained and tested either on regular images (left) or on images mapped on the retinopic space using a log-polar grid (right). Red predicts the presence of the label ’wolf’, blue its absence.
To quantify this effect, we tested the networks with the likelihood map protocol on an $8\times 8$ fixed grid of fixation points and an input sample with a 30% ratio see Table 1. Using the heat map extracted from the key points of the Animal 10k (Yu et al., 2021) dataset as ground truth, "in" represents coordinates inside an animal (and "out" coordinates outside an animal, see Appendix Figure 7). For each point in the $8\times 8$ grid, a likelihood value is obtained (probability of the presence of an animal). Next, we calculate the average likelihood for all points that are inside the zone corresponding to the animal (likelihood "in"), as well as the average likelihood for the zone that does not contain the animal (likelihood "out"). Next, we compare the values obtained in the "in" zone with those obtained in the "out" zone. A higher contrast indicates a better ability of the network to identify regions of interest in an image. For the VGG16, both performed well on the task, although the Cartesian tends to maintain a high likelihood outside the area of interest. For the ResNet101 networks, the Cartesian version of the network seems to perform much less well than the Retinotopic version in this exercise (see Table 1). If we consider a good categorization to be a high average probability on "in" coordinates (or a low probability on "out" coordinates), then in general networks using Retinotopic meshes tend to be more contrasted than networks using Cartesian meshes, which is more apparent in the case of ResNet101, see Table 1).
@@
Table 1: Likelihood maps results for the Vgg16 and ResNet101 networks and as computed on the Animal 10k (Yu et al., 2021) dataset. In: Average probability for a fixation point in a position where an animal is present. Out: Average probability for a fixation point outside the position where an animal is present. Ratio: Average probability in positions where there is an animal on the average probability outside positions where there is an animal (In/Out). We highlight for each network the mapping which reaches best performances for all three conditions.
Vgg16 ResNet101
Cartesian Retinotopic Cartesian Retinotopic
In 0.92 0.76 0.73 0.80
Out 0.77 0.49 0.69 0.54
Ratio 1.19 1.59 1.06 1.50
3.4 Task Dependence of Localization
@@
We note that the likelihood maps generated by the model will vary depending on the label or class being predicted for a given input image. Specifically, the spatial distribution of high likelihood regions in the map differs depending on the label the model is trying to identify, or a class of objects defined by a set of labels. This suggests that the model learns to focus on different discriminative regions for different classes. The different likelihood maps provide insight into how the model’s spatial attention changes according to the visual patterns that drive each classification decision. We tested this hypothesis on an example image by showing likelihood maps for the classes ’animal’, ’dog’, and ’cat’. As expected, the different regions are separated accordingly (see Figure 4).
@@
Figure 4: Likelihood maps generated by Vgg16 networks trained and tested on retinotopic space using the log-polar grid. Searching for the (A)  "animal", (B) "dog", or (C) "cat" label in an image.
@@
The observation that likelihood maps vary by predicted class provides an exciting opportunity for future work. By modeling visual search as a path through these conditional spatial distributions, we may gain insight into how human gaze is affected during visual tasks. Attentional mechanisms have long been shown to guide eye movements in humans (Yarbus, 1961). This work therefore presents promising opportunities to use deep networks as a means to unravel cognitive processes through computational modeling.
4 Discussion
4.1 Retina-Inspired Mapping Enhances the Robustness of CNNs
@@
A first and main result of this study is to demonstrate the excellent ability of standard deep CNNs to deal with retinotopic inputs, even though this mapping represents a profound transformation of the visual inputs. The ResNet101 and Vgg16 networks adapt easily to inputs where a large part of the image is highly compressed in the periphery and the spatial arrangement is significantly disturbed. The accuracy rates achieved with retinotopic inputs are equivalent to those of the original models.
@@
In addition, the log-polar transformation has the advantage of better invariance to zoom and rotation. However, this invariance comes at the cost of reduced invariance to translation. For images that are not centered on the region of interest, one would have to shift the fixation point to the region of interest, similar to eye saccades.
@@
The integration of a retinotopic mapping approach holds great promise for improving the efficiency and accuracy of image processing tasks. Our results are consistent with physiological data on ultrafast image categorization, which show that human accuracy in recognizing briefly flashed images of animals is robust to rotation (Fabre-Thorpe, 2011). The log-polar compression used in our approach allows a seamless extension to larger images without a significant increase in computational cost.
4.2 Relation to Pre-attentive Mechanisms
@@
In our case, it seems that the retinotopic mapping allows for a more precise localization of the category of interest, e.g., an animal, compared to off-the-shelf pre-trained networks using a Cartesian representation. It also gives us insight into the features on which our networks actually rely. Such information can be compared with physiological data (Crouzet, 2011), used to design better CNNs, and ultimately allow physiological tests to be proposed to further explore the features needed to classify a label of interest. In particular, by focusing on the point of fixation with the highest probability in the likelihood maps, we could consider refining the training of the network to our retinotopic mapping.
4.3 Introducing Eye Movements
Building on these observations, simulating human saccadic eye movement patterns during visual tasks provides an exciting opportunity to gain further insight into these mechanisms using such networks. A protocol that iteratively classifies image patches corresponding to foveated regions in a manner that mimics eye movements could reveal how network performance is spatially modulated across the visual field. Comparing classification accuracy under different saccade strategies, such as selecting the most uncertain or most likely location at each step, would provide valuable information about how network attention operates. This framework also allows the modeling of popular biological saccade strategies, allowing direct comparison with human visual search behavior.
Overall, implementing foveated classification with algorithmic saccades would provide a powerful method for validating existing attentional mechanisms in these networks, as well as inspiring new architectural innovations through embodied, task-driven visual attention modeling. This line of research should be particularly well fitted for a dual-pathway model which is well adapted to infer ego-motions (Bakhtiari et al., ????; Mineault2021). Finally, the implementation of this robust categorization, coupled with refined localization of a label of interest and optimal saccade selection, could allow us to extend this study to a more complex task. One such task is visual search (i.e., the simultaneous localization and detection of a visual target), and the likelihood maps could provide the underlying pre-attentive mechanisms on which its effectiveness seems to depend.
Appendix A Appendix
A.1 Robustness to rotations on the Animal 10k dataset
@@
Figure 5: Average accuracy over the Animal 10k (Yu et al., 2021) dataset, shown for both retrained and pre-trained networks with different input image rotations. The rotation is applied around the fixation point with an angle ranging from $-180^{\circ}$ to $+180^{\circ}$ (in steps of $15^{\circ}$). We tested each network (Vgg16 or ResNet101) either with raw images in Cartesian coordinates or with retinotopic mapping. The dotted line represents chance level.
A.2 Creating Multiple Views for Likelihood Maps
@@
Figure 6: Generating different views of a single image to compute likelihood maps. (A) For the networks with Cartesian inputs, we used a regular grid of $8\times 8$ fixation points, which allows samples to be cropped, highlighting a particular view. As shown in (B), this produces a batch of images that can be used to generate likelihood maps. (C)) We used a similar grid to generate batches of retinotopic inputs, as shown in (D)). In (B) & (D) each sample corresponds to $30\%$ of the input.
A.3 Ground Truth Heat Maps
Figure 7: (A) The original image of the Animal 10k dataset. (B) A heat map constructed by fitting Gaussians to the key points of the Animal 10k dataset. (C) The heat map constructed in (B) is normalized and reduced to a resolution of $8\times 8$ to be used as ground truth when evaluating the heat map. A threshold (0.2) is applied to reduce the heat map field to the assumed contour of the animal.
@@ A.3 Ground Truth Heat Maps
Figure 8: Likelihood maps computed on a representative image using $8\times 8$ points of fixation, (as computed on the whole Animal 10k (Yu et al., 2021) dataset), (top) with a Vgg16, or (bottom) a ResNet101) network, trained and tested on either regular images (left) or images mapped in retinopic space using a log-polar grid (right). Red predicts the presence of the label ’animal’, blue its absence.

### v2

@@ Abstract
We fine-tuned the pre-trained ResNet networks on the ImageNet dataset with retinotopic transformed images then compare the performance of the networks trained on the original dataset (with a Cartesian mapping) with the performance of our retrained network on the transformed dataset defined in retinotopic space. The results shows that the networks retrained on the transformed images have similar categorization accuracy as the network pre-trained on regular images (see Figure 2-A & B). This is surprising given that the networks used in this retraining process were previously trained on regular images and that images with a log-polar transformation show a high degree of distortion, in particular a compression of visual information around the fixation point and a degradation of textures in the periphery, see Figure 1. One hypothesis is that the degradation of texture during the frame of reference change may cause the network to rely on shape rather than texture [da_costa_convolutional_2024].
Robustness of CNNs
Deep learning algorithms have made tremendous progress in the last few years. For some visual recognition tasks, such as the ImageNet challenge [31], they have achieved superhuman accuracy. However, one limitation that remains is their vulnerability to adversarial attacks. Studies have shown that these learned models can be fooled by imperceptible modifications to images that are undetectable to humans [38]. Yet these small distortions cause the algorithms to incorrectly classify examples with high confidence [39]. This vulnerability makes deep networks unstable and unsuitable for use in safety-critical domains like medicine, autonomous vehicles, or other life-or-death situations. Before deep learning can be relied upon for such applications, researchers must find ways to make these models more resilient to adversarial examples and introduce human-level robustness to ensure that mistakes do not have dangerous consequences in the real world.
Attacking classical CNNs with a geometrical rotation
More specifically, our approach consists first to choose the rotation angle that maximizes the loss, then to infer the most probable label for that particular angle. As a result, we can compute the concordance between the predicted label $\hat{y}$ for the image rotated at the angle $\bar{\theta}$ with the worst loss with respect to the ground truth label $y$. Using this procedure, we calculated the overall accuracy on the entire test set, quantifying the network’s brittleness to natural image rotations.
Our experiments showed that while ResNet101 reached nominally 74% baseline accuracy, rotation attacks significantly reduced this performance. The application of the maximally deceptive rotation to each image reduced the accuracy to 39%. These results were confirmed by performing the same procedure on the Animal 10k [32] dataset (see Appendix Figure 7). In addition, we computed for any given rotation the average accuracy over the test dataset, showing that this average accuracy dropped sharply compared to baseline accuracy, confirming that this simple geometrical transformation mislead the networks. The decline was steepest between two cardinal rotations, demonstrating limited rotational invariance compared to humans [40].
Fig 2: To illustrate the brittleness of classical CNNS, we show the effect of rotation-based attack by plotting accuracy over a sample of $1000$ images from the Imagenet [31] dataset, for (A) For all the networks, we plot the accuracy over the raw data set (without rotation, in light color), or on the image rotated at the angle $\bar{\theta}$ with the worst loss (rotation-based attack, in full color). (B) Average accuracy over a sample of 1500 images from the Imagenet [31] dataset, shown for both retrained and pre-trained networks with different input image rotations. The rotation is applied around the fixation point with an angle ranging from $-180^{\circ}$ to $+180^{\circ}$ (in steps of $15^{\circ}$).
We test each network for both frameworks (Cartesian and Retinotopic). We observe a symmetry with respect to the horizontal flips and only show the positive angles here. This shows that the Cartesian network has degraded performance compared to ResNet101, and notably, rotating images may have an adversarial effect on categorization performance, an effect that is less observed for the retinotopic network.
@@
Notably, these results show that while the Cartesian network has some degradation for different rotations, the categorization results are much more invariant for the network that includes a retinotopic mapping (see Figure 2-C). This phenomenon is a consequence of the translation invariance imposed by the structure of the CNNs. When applied to retinotopic mapping, this translation invariance in retinotopic space translates into rotation and zoom invariance in visual space [19].
Visual object localization : Protocol
@@
Deep convolutional networks output a vector of real numbers, often interpreted as the log-probability on the label space (“logit”), as implemented through cross-entropy loss. Applying the softmax function allows the output to be interpreted as a probability vector, assigning a probability of presence to each of the $1000$ labels (likelihood score). This allows us to make a binary decision (“present” or “not present”), e.g. by selecting the label corresponding to the highest likelihood. In our setting, we can also take different views of a large image and compute the likelihood for each of them, allowing us to compare which view provides the best likelihood. Views can consist, for example, of focusing on regions of the image centered on different fixation points, with the fixation points aligned on a regular grid in visual space.
@@
This sample is then transformed or not by the retinotopic mapping before being used as input for the corresponding network, see Appendix Figure 8-B & D. Conveniently, a collection of samples for different fixation points can be processed as a single batch. This protocol defines a likelihood map for any given network as the likelihood of categorizing the presence of a label of interest inferred at regularly spaced fixation points in the image. As can be observed qualitatively in the example image, the Cartesian representation induces uniform activation across the whole image, while the retinotopic mapping induces differential activation inside and outside the object of interest.
Fig 3: Likelihood maps computed on a representative image using $11\times 1$ 11 points of fixation with networks, trained and tested either on regular images (Left) or on images mapped on the retinopic space using a log-polar grid (Right).
Visual object localization : Quantification
We sub sampled each image with a grid of equidistant viewpoints at a resolution of $11\times 1$ 11. At each viewpoint, the largest possible sample is cropped. Thus a minimum sample with a 30% ratio at the border and the whole image at the center. From the Cartesian or retinotopic reference frame, this sample is then resized to a $224\times 224$ resolution to match the optimal size for the CNN before processing. We selected some key metrics to compare the retinotopic or Cartesian referential. An indicator of the correct map activation position is the Energy-Based Pointing Game accuracy [41] where the localization is a success when the peak activation of the heat map of a given label is located inside the ground truth bounding box. In addition, we chose to track the ratio of the activation inside versus outside the box: the higher the ratio and the greater is the contrast of the heat map.
Using the heat map extracted from the key points of the Animal 10k [32] dataset as ground truth, “in” represents coordinates inside an animal (and “out” coordinates outside an animal, see Appendix Figure 9). For each point in the $11\times 1$ 11 grid, a likelihood value is obtained (probability of the presence of an animal). Next, we calculate the average likelihood for all points that are inside the zone corresponding to the animal (likelihood “in”), as well as the average likelihood for the zone that does not contain the animal (likelihood “out”). Next, we compare the values obtained in the “in” zone with those obtained in the “out” zone. A higher contrast indicates a better ability of the network to identify regions of interest in an image. If we consider a good categorization to be a high average probability on “in” coordinates (or a low probability on “out” coordinates), then in general networks using Retinotopic meshes tend to be more contrasted than networks using Cartesian meshes, which is more apparent in the case of ResNet101, see Table 1).
@@
Table 1: Likelihood maps results for the ResNet networks and as computed on the Imagenet [31] dataset. In: Average probability for a fixation point in a position where an animal is present. Out: Average probability for a fixation point outside the position where an animal is present. Ratio: Average probability in positions where there is an animal on the average probability outside positions where there is an animal (In/Out). PG: In this context, the term “PG” is used to denote the degree of accuracy achieved in the Pointing Game. IoU-Peak: Represent the maximum value of the Iou reached by the network. IoU-T: In this context, “IoU-T” refers to the IoU Threshold, that is to say, the value at which the IoU reaches zero for each network. We highlight the mapping which reaches the best performances for all three conditions for each network.
Resnet18 Resnet50 ResNet101
Cartesian Retinotopic Cartesian Retinotopic Cartesian Retinotopic
In 0.46 0.31 0.53 0.59 0.47 0.63
Out 0.17 0.09 0.40 0.40 0.38 0.45
Ratio +38% +27% +33% +44% +23% +41%
PG 72% 80% 64% 74% 66% 76%
IoU-Peak 0.29 0.29 0.28 0.29 0.28 0.29
IoU-T 0.9 0.99 0.78 0.88 0.82 0.95
Accuracy after “saccades” protocol
Fig 5: Sample saccade to maximum likelihood position
In this part of the study, we focused on finding a label of interest in including a large number of fixation points per image. We applied this protocol to the 50,000 images in the validation set of the ImageNet data set. If we only stop at the best position (Top 1), the performance of the networks is degraded compared to their accuracy without saccades, and the same is true for Top 5 (compared to the performance of Top 5 without saccades, not shown here). Here, a saccade correspond to the selection of another fixation point in the grid. In this part of the study we are trying to explore if there is a point of fixation in the visual scene that could increase the categorisation of the networks. By selecting the fixation point (saccade), either the most salient to the target label (’priors’) or the most salient independent to the target label (’no priors’). The results indicate that networks using the retinotopic reference frame appear to maximize prediction after moving the fixation point, as the mean accuracy is better in this test. Note that since the networks are trained on the boxes, their pre-saccadic acuity is a test in itself, as they are presented with the full image. Although cartesian networks perform better in validation during training, they lose their advantage on the full image. Those results could reflext a smaller proportion of out of distribution categorisation for the Retinotopic framework due to its enhance robustness to current transformation such as rotation and zoom.
Task dependence of localization
@@
Another dimension that could be usefull in the definition of the likelihood map is the dimension of the label. In fact, the networks perform a discriminative categorisation among 1000 Imagenet labels. We observe that the likelihood maps generated by the model will vary depending on the label or class being predicted for a given input image. Specifically, the spatial distribution of high likelihood regions in the map differs depending on the label the model is trying to identify, or a class of objects defined by a set of labels. This suggests that the model learns to focus on different discriminative regions for different classes. The different likelihood maps provide insights into how the model’s spatial attention changes according to the visual patterns that drive each classification decision. We tested this hypothesis on an example image by showing likelihood maps for the classes “animal”, “dog”, and “cat”. As expected, the different regions are separated accordingly (see Figure 6).
@@
Fig 6: Likelihood maps generated by Resnet101 networks trained and tested on retinotopic space using the log-polar grid. Searching for the (A)  “animal”, (B) “dog”, or (C) “cat” label in an image.
@@
The observation that likelihood maps vary by predicted class provides an exciting opportunity for future work. By modeling visual search as a path through these conditional spatial distributions, we may gain insight into how human gaze is affected during visual tasks. Attentional mechanisms have long been shown to guide eye movements in humans [4]. This work therefore presents promising opportunities to use deep networks as a means to unravel cognitive processes through computational modeling.
Discussion
Retina-inspired mapping enhances the robustness of CNNs
@@ Discussion
A first and main result of this study is to demonstrate the excellent ability of standard deep CNNs to deal with retinotopic inputs, even though this mapping represents a profound transformation of the visual inputs. The ResNet networks adapt easily to inputs where a large part of the image is highly compressed in the periphery and the spatial arrangement is significantly disturbed. The accuracy rates achieved with retinotopic inputs are equivalent to those of the original models.
@@
In addition, the log-polar transformation has the advantage of better invariance to zoom and rotation. However, this invariance comes at the cost of reduced invariance to translations. For images that are not centered on the region of interest, one would have to shift the fixation point to the region of interest, similar to eye saccades.
@@
The integration of a retinotopic mapping approach holds great promise for improving the efficiency and accuracy of image processing tasks. Our results are consistent with physiological data on ultrafast image categorization, which show that human accuracy in recognizing briefly flashed images of animals is robust to rotation [42]. The log-polar compression used in our approach allows a seamless extension to larger images without a significant increase in computational cost.
Relation to Pre-attentive Mechanisms
@@
In our case, it seems that the retinotopic mapping allows for a more precise localization of the category of interest, e.g., an animal, compared to off-the-shelf pre-trained networks using a Cartesian representation. It also gives us insight into the features on which our networks actually rely. Such information can be compared with physiological data [43], used to design better CNNs, and ultimately allow physiological tests to be proposed to further explore the features needed to classify a label of interest. In particular, by focusing on the point of fixation with the highest probability in the likelihood maps, we could consider refining the training of the network to our retinotopic mapping.
Introducing eye movements
Building on these observations, simulating human saccadic eye movement patterns during visual tasks provides an exciting opportunity to gain further insights into these mechanisms using such networks. A protocol that iteratively classifies image patches corresponding to foveated regions in a manner that mimics eye movements could reveal how network performance is spatially modulated across the visual field. Comparing classification accuracy under different saccade strategies, such as selecting the most uncertain or most likely location at each step, would provide valuable information about how network attention operates. This framework also allows the modeling of popular biological saccade strategies, allowing direct comparison with human visual search behavior.
Overall, implementing foveated classification with algorithmic saccades would provide a powerful method for validating existing attentional mechanisms in these networks, as well as inspiring new architectural innovations through embodied, task-driven visual attention modeling. This line of research should be particularly well-fitted for a dual-pathway model which is well adapted to infer ego-motions [44, 45]. Finally, the implementation of this robust categorization, coupled with refined localization of a label of interest and optimal saccade selection, could allow us to extend this study to a more complex task. One such task is visual search (i.e., the simultaneous localization and detection of a visual target), and the likelihood maps could provide the underlying pre-attentive mechanisms on which its effectiveness seems to depend. For more information, see LABEL:S1_Appendix.
Acknowledgments
Authors received funding from the ANR project number ANR-20-CE23-0021 (”AgileNeuroBot”) and from the french government under the France 2030 investment plan, as part of the Initiative d’Excellence d’Aix-Marseille Université – A*MIDEX grant number AMX-21-RID-025 “Polychronies”. This work was granted access to the HPC resources of Aix-Marseille Université financed by the project Equip@Meso (ANR-10-EQPX-29-01) of the program “Investissements d’Avenir” supervised by the Agence Nationale de la Recherche. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
Appendix
Robustness to rotations on the Animal 10k dataset
@@ Acknowledgments
Fig 7: Average accuracy over the Animal 10k [32] dataset, shown for both retrained and pre-trained networks with different input image rotations. The rotation is applied around the fixation point with an angle ranging from $-180^{\circ}$ to $+180^{\circ}$ (in steps of $15^{\circ}$). We tested each network either with raw images in Cartesian coordinates or with retinotopic mapping. The dotted line represents chance level.
Creating Multiple Views for Likelihood Maps
@@
Fig 8: Generating different views of a single image to compute likelihood maps. (A) For the networks with Cartesian inputs, we used a regular grid of $8\times 8$ fixation points, which allows samples to be cropped, highlighting a particular view. As shown in (B), this produces a batch of images that can be used to generate likelihood maps. (C)) We used a similar grid to generate batches of retinotopic inputs, as shown in (D)). In (B) & (D) each sample corresponds to $30\%$ of the input.
Ground Truth Heat Maps
Fig 9: (A) The original image of the Animal 10k dataset. (B) A heat map constructed by fitting Gaussians to the key points of the Animal 10k dataset. (C) The heat map constructed in (B) is normalized and reduced to a resolution of $8\times 8$ to be used as ground truth when evaluating the heat map. A threshold (0.2) is applied to reduce the heat map field to the assumed contour of the animal.
@@
Fig 10: Likelihood maps computed on a representative image using $8\times 8$ points of fixation, (as computed on the whole Animal 10k [32] dataset), (top) with a Vgg16, or (bottom) a ResNet101) network, trained and tested on either regular images (left) or images mapped in retinopic space using a log-polar grid (right). Red predicts the presence of the label ’animal’, blue its absence.